C) 11 of (The 1 of1> Znd gear Estmate the average acceleration of the car in third gear f (s) 10 20 30 40
Answer #1

Part A -

First gear motion -

v1 = 0

v2 = 14 m/s

Time taken, t1 = 4 s

Average acceleration of the car in first gear, a1 = Change in velocity / time taken

= (v2 - v1) / t1 = (14 - 0) / 4 = 3.5 m/s^2

Part B -

Like above -

v3 = 24 m/s

v4 = 36 m/s

t3 = 16 - 12 = 4 s

Therefore, average acceleration of the car in third gear, a3 = (v4 - v3) / t3 = (36 - 24) / 4 = 3.0 m/s^2